Find the buoyant force exerted on a ball with radius of 2 meters, when the ball is entirely immersed in the water.
Answer #1

here,

radius, r = 2 m

density of water, p = 1000 Kg/m^3

buoyant force on the ball = weight of the water displaced by the ball

buoyant force on the ball = p * V * g

buoyant force on the ball = 1000 * 4pi/3 * 2^3 * 9.8

buoyant force on the ball = 3.284 *10^5 N

the buoyant force on the ball is 3.284 *105 N